OpenStudy (anonymous):

find the points on the curve y=4x^3 -6x^2- 24x +30 where the tangent is parallel to the x axis. please help and show all steps!

OpenStudy (anonymous):

parallel to x-axis mean gradient = 0 could be maximun or mininum point find its dy/dx, is 12x²-12x-24 0= 12x²-12x-24 =x²-x-2 =(x-2)(x-1) that mean when x=2 or 1, the gradient is 0, to find the y value, just input x's value into the origibnal equation.